OpenStudy (anonymous):

y = (3x + 2)^4(2x - 5)^2 how do i find derivative

sam (.sam.):

Product rule + Chain rule \[(3 x+2)^4 \left(\frac{d}{dx}\left((2 x-5)^2\right)\right)+(2 x-5)^2 \left(\frac{d}{dx}\left((3 x+2)^4\right)\right)\] \[4 (2 x-5) (3 x+2)^4+12 (2 x-5)^2 (3 x+2)^3\] \[4 (2 x-5) (3 x+2)^3 (9 x-13)\]

OpenStudy (anonymous):

where did the 4 come from in the beginning

OpenStudy (anonymous):

Derivative of x^n = n x^ ( n -1)

OpenStudy (anonymous):

ok so how did u get from step 2 to step 3

OpenStudy (anonymous):

factorization!